OpenStudy (jgirl128):

PLEASE help?? The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work. I don't just want the answer, I want someone to help me work it out. Thanks!!

OpenStudy (anonymous):

Given: \[\huge d = 0.5g t^2 \]

OpenStudy (shamim):

g=d/(0.5*t^2)

OpenStudy (anonymous):

\[\huge g = \frac{d}{0.5t^2}\]

OpenStudy (jgirl128):

how did you get that?

OpenStudy (jgirl128):

the g=d/(0.5*t^2), I mean

OpenStudy (anonymous):

Dividing both side by 0.5 t^2

OpenStudy (anonymous):

Now substitute d=490 and t 10 in the above formula

OpenStudy (jgirl128):

so it would be d 0.5t _____ = ______ 0.5t2 0.5t2

OpenStudy (jgirl128):

and then substitute d over 0.5t2 for 490 over 10?

OpenStudy (jgirl128):

or is that wrong

OpenStudy (anonymous):

\[\huge g= \frac{d}{0.5t^2}= \frac{490}{0.5 \times 10^2}\] \[\huge g= \frac{490}{0.5 \times 100}= \frac{490}{50}= \frac{49}{5}\] \[\huge g= \frac{49}{5} =9.8\] hence g= 9.8 feet/sec^2

OpenStudy (jgirl128):

so the acceleration would be 9.8 feet/sec²?

OpenStudy (anonymous):

Yes

OpenStudy (jgirl128):

anyways, thanks for helping!! :)